OpenStudy (anonymous):

Hi. I read from your past questions answered that you have solutions manual for integral and differential calculus by feliciano and uy? Can you help me find? I can't access your link. Please help.

OpenStudy (zzr0ck3r):

who are you talking to ?

OpenStudy (zzr0ck3r):

next time I would just message that person. As pirating books on here is not the best idea....

OpenStudy (anonymous):

Intended for Machida but, Anyone would be great, tho. :) Ops sorry.

OpenStudy (anonymous):

Find the equations of the tangent and normal to the graph of the given function at the given point. 1. y+3x^2 - 2x + 1, (2,9) 2. y=1=3 square root of x, (4,7) 3.y = x square root of x - 1 , (5,10) 4. y^2 = x^3 over 4-x , (2,2) 5. y = 2 over x , (1,2)